# Shijian Li1,2, Tao Li4,5,1, Gang Pan1,2,5 1

### v1

@@ Appendix G Architecture Comparison on Our Pre-training Datasets
In this paper, we use different model architecture and pre-training dataset compared to BIOT (Yang et al., 2023) and LaBraM (Jiang et al., 2024). To further evaluate whether the model architecture or the pre-training dataset contributes more to performance improvement, we conduct a comparison experiment. Specifically, we pre-trained the architecture of BIOT and LaBraM on our pre-training dataset with the same settings as CBraMod, and compared the performance with BIOT (original), LaBraM (original) and CBraMod on downstream datasets. BIOT (ours) and BIOT (original) exhibit very similar performance. LaBraM (ours) generally achieves slightly better performance than LaBraM (original), particularly on the FACED dataset. But CBraMod performs significantly better than LaBraM (ours) on all datasets. All results empirically prove that the model architecture contributes more to the performance improvement compared to the larger size of pretraining dataset.
@@ Appendix I Ablation Study on Time-Domain and Frequency-Domain Signals
Appendix Q Discussion
Q.1 Implication
Q.2 Limitation
Q.3 Future Work
Appendix R A leave-one-subject-out Comparison with EEG-SimpleConv on BCIC-IV-2a
@@
Appendix S Comparison on Convergence Speed with Supervised Model

### v2

@@ Appendix G Architecture Comparison on Our Pre-training Datasets
In this paper, we use different model architecture and pre-training dataset compared to BIOT (Yang et al., 2023) and LaBraM (Jiang et al., 2024). To further evaluate whether the model architecture or the pre-training dataset contributes more to performance improvement, we conduct a comparison experiment. Specifically, we pre-trained the architecture of BIOT and LaBraM on our pre-training dataset with the same settings as CBraMod, and compared the performance with BIOT (original), LaBraM (original) and CBraMod on downstream datasets. The experimental results are shown in Figure 8.
BIOT (ours) and BIOT (original) exhibit very similar performance. LaBraM (ours) generally achieves slightly better performance than LaBraM (original), particularly on the FACED dataset. But CBraMod performs significantly better than LaBraM (ours) on all datasets. All results empirically prove that the model architecture contributes more to the performance improvement compared to the larger size of pretraining dataset.
@@ Appendix I Ablation Study on Time-Domain and Frequency-Domain Signals
Appendix Q A leave-one-subject-out Comparison with EEG-SimpleConv on BCIC-IV-2a
@@
Appendix R Comparison on Convergence Speed with Supervised Model
@@
Appendix S Discussion
S.1 Implication
S.2 Limitation
S.3 Future Work